ganeshie8 (ganeshie8):

\((2n)! + 1 = (2n+1)^k\) find all solutions over natural numbers

ganeshie8 (ganeshie8):

this is a follow up of earlier problem http://openstudy.com/study#/updates/566af694e4b090814d904aea

OpenStudy (bubblegum.):

\((2n+1)^k=\left(\begin{matrix}k\\ 0\end{matrix}\right)(2n)^01^k+\left(\begin{matrix}k\\ 1\end{matrix}\right)(2n)^11^{k-1}+....\left(\begin{matrix}k \\ n\end{matrix}\right)(2n)^k1^{k-1}\) \((2n+1)^k=1+\left(\begin{matrix}k\\ 1\end{matrix}\right)(2n)^11^{k-1}+....\left(\begin{matrix}k \\ k\end{matrix}\right)(2n)^k1^{k-1}\) \((2n+1)^k=1+k\left( ^{k-1}C_0+ ^{k-1}C_1....+^{k-1}C_{k-1} \right)\) \((2n+1)^k=1+k2^{k-1}\) now we just put in our equation \((2n)!+1=(2n+1)^k\) \((2n)!+1=1+k2^{k-1}\) \((2n)!=k2^{k-1}\) now we hav to find solutions to this equation :/

OpenStudy (bubblegum.):

in the 1st line the coefficient of rightmost term -> \( \left(\begin{matrix}k \\ k\end{matrix}\right)\)

OpenStudy (bubblegum.):

seems like i did something wrong n=0, k=1 satisfies the equation -> \((2n)!=k2^{k-1}\) but it doesn't satisfy \((2n)!+1=(2n+1)^k\)

ganeshie8 (ganeshie8):

Except for that mistake, it is a great start ! Here I have fixed the mistake : \((2n)!+1 = (2n+1)^k \) \(\implies (2n)! = (2n+1)^k-1^k\) factor right hand side using the formula http://math.stackexchange.com/questions/117660/proving-xn-yn-x-yxn-1-xn-2-y-x-yn-2-yn-1 \(\implies (2n)! = (2n)\sum\limits_{i=0}^{k-1}(2n+1)^i\) \(\implies (2n-1)! = \sum\limits_{i=0}^{k-1}(2n+1)^i\)

ganeshie8 (ganeshie8):

let \(2n+1 = m\), above equation is same as : \(\implies (m-2)! = \sum\limits_{i=0}^{k-1}m^i \)

ganeshie8 (ganeshie8):

Next we may use the earlier proven result : \((m-2)!\) is divisible by \(m-1\)

Parth (parthkohli):

so this means we only have a limited number of solutions, then? n = 1, k = 1 is a pretty trivial one

ganeshie8 (ganeshie8):

Yes, there are no solutions for \(n\gt 2\)

OpenStudy (bubblegum.):

yeah n=2,k=2 is also a sol

ganeshie8 (ganeshie8):

Yes thats all... I cooked up the problem from this : http://math.stackexchange.com/questions/1195342/when-does-p-1-1-pk-hold

ganeshie8 (ganeshie8):

that link has a simpler version of the same problem...

Parth (parthkohli):

oh now we have to prove the rest :|

Parth (parthkohli):

it's weird though... the fact that we can freely choose our \(k\) makes this unbelievable

ganeshie8 (ganeshie8):

let \(2n+1 = m\), above equation is same as : \(\implies (m-2)! = \sum\limits_{i=0}^{k-1}m^i \) If above equation holds, then it also holds under mod \(m-1\) : \(\implies 0\equiv \sum\limits_{i=0}^{k-1}1^i\pmod{m-1} \) \(\implies 0\equiv k \pmod{m-1} \) \(\implies k=t(m-1) \)

ganeshie8 (ganeshie8):

so above shows that we cannot choose \(k\) freely...

Parth (parthkohli):

yes of course, we can boil the restrictions down but I get a similar feeling for this like I do for the FLT.

Parth (parthkohli):

of course the proof must be simpler for this, but for a newbie they're all the same :)

ganeshie8 (ganeshie8):

proof of FLT is very easy for you

ganeshie8 (ganeshie8):

i bet you can prove it on ur own if you try

Parth (parthkohli):

loooooool

Parth (parthkohli):

no I mean Fermat's Last Theorem

ganeshie8 (ganeshie8):

i thought you're talking about little fermat

Parth (parthkohli):

sorry I forgot they used FLT for little

ganeshie8 (ganeshie8):

FLT is beyong the scope of openstudy

ganeshie8 (ganeshie8):

*beyond

Parth (parthkohli):

openstudy? maybe even stackexchange

ganeshie8 (ganeshie8):

i think anyone who took abstract algebra can understand the wiles proof

Parth (parthkohli):

were you able to understand it?

ganeshie8 (ganeshie8):

https://www.math.wisc.edu/~boston/869.pdf

ganeshie8 (ganeshie8):

i could never get past the first few pages

ganeshie8 (ganeshie8):

This book will describe the recent proof of Fermat’s Last Theorem by Andrew Wiles, aided by Richard Taylor, for graduate students and faculty with a reasonably broad background in algebra. It is hard to give precise prerequisites but a first course in graduate algebra, covering basic groups, rings, and fields together with a passing acquaintance with number rings and varieties should suffice.

Parth (parthkohli):

it's weird to think how much Galois had done by the time he was my age :(

ganeshie8 (ganeshie8):

Galois and Abel contributed a lot to algebra and both died young, so unfortunate

Parth (parthkohli):

While we're at them, Ramanujan too.

Parth (parthkohli):

though his contribution can only be appreciated by his fellow mathematicians.

Parth (parthkohli):

ok how do you prove this lol

ganeshie8 (ganeshie8):

It follows \(k = t(2n)\) for \(n\gt 2\), substituting this in the given equation : \[(2n)! + 1 = (2n+1)^{t(2n)}\]

ganeshie8 (ganeshie8):

Above equation is impossible because the left hand side is always less than the right hand side

ganeshie8 (ganeshie8):

\[(2n)!+1 \lt (2n+1)(2n+1)\cdots (2n\text{ times}) = (2n+1)^{2n}\]

ganeshie8 (ganeshie8):

that proves no solutions for \(n > 2\) just need to check the cases \(n=1,2\) manually

OpenStudy (bubblegum.):

nice :)

ganeshie8 (ganeshie8):

the entire proof hinges on the result from previous problem : \((n-2)!\) is divisible by \(n-1\) whenever \(n\ne p+1\) and \(n\gt 5\)

ganeshie8 (ganeshie8):

only because of that result is it possible to simplify the problem by considering mod \(m-1\)

ganeshie8 (ganeshie8):

I do not see any other way to work this problem If we are not familiar with that result, I think we will be stuck forever..

OpenStudy (bubblegum.):

can we differentiate factorials? will that be easy and look good? like \(\large d \frac{ (n)! }{ d(n) }\)

ganeshie8 (ganeshie8):

we may use the continuous version of factorial, the gamma function

ganeshie8 (ganeshie8):

\[\Gamma(t) = \int\limits_{0}^{\infty}x^{t-1}e^{-x}\,dx\]

ganeshie8 (ganeshie8):

This is an extension of the regular factorial defined for all complex numbers except nonnegatice integers

ganeshie8 (ganeshie8):

\(\Gamma(t+1) = t\Gamma(t)\)